# Stronger Together: On-Policy Reinforcement Learning for Collaborative LLMs

### v2

@@ A.1.1 Math reward design
$c^{\mathrm{Tool}}_{\mathrm{fmt}}=0.10,\quad c^{\mathrm{Tool}}_{\mathrm{tool}}=0.10,\quad c^{\mathrm{Tool}}_{\mathrm{step}}=0.80.$
@@
$s^{\mathrm{Tool}}_{\mathrm{step},k}=\begin{cases}\textsc{NumEq}_{\delta}\!\big(\tilde{y}_{k},\,y^{\star}\big),&\text{if execution emits numeric }\tilde{y}_{k}\text{ at }k,\\[2.0pt] 0,&\text{otherwise},\end{cases}$
$m^{\mathrm{Tool}}_{k}=\mathbf{1}\{y^{\star}\ \text{and logs available (}\textsc{MathVerifier}\text{)}\}.$
A.1.2 Code reward design
We consider code synthesis with unit tests at each turn $k$. Let $\mathcal{S}_{k}$ be the active test suite and
$p_{k}=\frac{1}{|\mathcal{S}_{k}|}\sum_{t\in\mathcal{S}_{k}}\mathbf{1}\{\textsc{Run}(t,\text{code}_{k})=\textsf{pass}\}\in[0,1].$
Team reward is dense:
$r^{\mathrm{team}}_{k}\;=\;p_{k}.$
Local rewards use fixed coefficients over verifiable sub-scores $s_{\ell,k}^{i}\in[0,1]$:
$r^{i,\mathrm{loc}}_{k}\;=\;\sum_{\ell}c^{i}_{\ell}\,s^{i}_{\ell,k},\qquad\sum_{\ell}c^{i}_{\ell}=1.$
Coder local design.
Coefficients (fixed):
Component scores:
$s^{\mathrm{Coder}}_{\mathrm{build},k}=\mathbf{1}\{\text{compiles/imports without syntax errors at }k\},$
$s^{\mathrm{Coder}}_{\mathrm{run},k}=\mathbf{1}\{\text{smoke subset runs without uncaught exceptions/timeout at }k\},$
$s^{\mathrm{Coder}}_{\mathrm{nr},k}=\begin{cases}\frac{1}{|G_{k-1}|}\sum_{t\in G_{k-1}}\mathbf{1}\{\textsc{Run}(t,\text{code}_{k})=\textsf{pass}\},&|G_{k-1}|>0,\\ 1,&|G_{k-1}|=0,\end{cases}$
where $G_{k-1}=\{t\in\mathcal{S}_{k-1}:\textsc{Run}(t,\text{code}_{k-1})=\textsf{pass}\}$. Mask:
$m^{\mathrm{Coder}}_{k}=\mathbf{1}\{\text{build/run logs and test diffs available at }k\}.$
Tester local design.
Coefficients (fixed):
$c^{\mathrm{Tester}}_{\mathrm{valid}}=0.20,\quad c^{\mathrm{Tester}}_{\mathrm{cov}}=0.80.$
Component scores:
@@ A.1.2 Code reward design
$s^{\mathrm{Tester}}_{\mathrm{cov},k}=\begin{cases}\min\!\left(1,\ \dfrac{\mathrm{MutScore}_{k}}{\tau_{\mathrm{mut}}}\right),&\text{mutation analysis available},\\[6.0pt] 0,&\text{otherwise},\end{cases}$
where $(x)_{+}=\max(x,0)$, $\mathrm{MutScore}_{k}\in[0,1]$ is the mutation score on golden code, $\mathrm{BrCov}_{k}\in[0,1]$ is branch coverage, and thresholds are fixed as
$\tau_{\mathrm{mut}}=0.60,\qquad\tau_{\mathrm{cov}}=0.10.$
Mask:
$m^{\mathrm{Tester}}_{k}=\mathbf{1}\{\text{test runner and mutation/coverage reports available at }k\}.$
A.1.3 Sudoku reward design
We consider $N{\times}N$ Sudoku with horizon $K$. Let $h_{k}$ be the dialogue/tool history at turn $k$ and $\textsc{Solved}(\cdot)$ check row/column/subgrid validity. Team reward is a sparse success signal at termination:
$r^{\mathrm{team}}_{k}\;=\;\mathbf{1}\{k{=}T,\ \textsc{Solved}(h_{k}){=}\textsf{true}\}\in\{0,1\}.$
We set the team–local mixing coefficient to a fixed number
$\lambda_{\mathrm{sudoku}}=0.60.$
For each agent $i\in\{\mathrm{Reasoner},\mathrm{Tool}\}$ at turn $k$, with verifiability mask $m^{i}_{k}\in\{0,1\}$, the per-agent learning reward is
$r^{i}_{k}\;=\;\lambda_{\mathrm{sudoku}}\,r^{\mathrm{team}}_{k}\;+\;(1-\lambda_{\mathrm{sudoku}})\,m^{i}_{k}\,r^{i,\mathrm{loc}}_{k}.$
Local rewards are convex combinations of component scores $s^{i}_{\ell,k}\in[0,1]$ with fixed coefficients $\{c^{i}_{\ell}\}$ summing to $1$:
$r^{i,\mathrm{loc}}_{k}\;=\;\sum_{\ell}c^{i}_{\ell}\,s^{i}_{\ell,k},\qquad\sum_{\ell}c^{i}_{\ell}=1.$
Reasoner local design.
Coefficients (fixed):
$c^{\mathrm{Reasoner}}_{\mathrm{fmt}}=0.15,\quad c^{\mathrm{Reasoner}}_{\mathrm{legal}}=0.55,\quad c^{\mathrm{Reasoner}}_{\mathrm{prog}}=0.30.$
Component scores at turn $k$ (let $G_{k}$ be the current grid, $G_{k-1}$ the previous grid; 0 denotes empty):
[…]
$s^{\mathrm{Reasoner}}_{\mathrm{legal},k}=\mathbf{1}\{\text{no row/column/subgrid duplicates in }G_{k}\},$
$s^{\mathrm{Reasoner}}_{\mathrm{prog},k}=\frac{1}{N^{2}}\sum_{r,c}\mathbf{1}\{G_{k-1}[r,c]{=}0,\ G_{k}[r,c]{\neq}0\}.$
@@ A.1.3 Sudoku reward design
$c^{\mathrm{Tool}}_{\mathrm{fmt}}=0.10,\quad c^{\mathrm{Tool}}_{\mathrm{exec}}=0.20,\quad c^{\mathrm{Tool}}_{\mathrm{san}}=0.70.$
@@
$s^{\mathrm{Tool}}_{\mathrm{fmt},k}=\mathbf{1}\{\text{API/schema valid; values in }[1,N];\ \text{indices in bounds}\},$
$s^{\mathrm{Tool}}_{\mathrm{exec},k}=\mathbf{1}\{\text{no runtime error/timeout when applying edits}\},$
$s^{\mathrm{Tool}}_{\mathrm{san},k}=\begin{cases}1,&\text{if all applied edits satisfy local Sudoku constraints},\\ 0,&\text{otherwise}.\end{cases}$
@@
A.1.4 Plan-Path reward design
We consider 2D grid path planning on a $H{\times}W$ map with horizon $K$ and four-neighborhood moves. Let $d_{k}$ be the Manhattan distance from the current position to the goal at turn $k$ and $d_{0}=\max(1,\text{initial distance})$ for normalization. Team reward is dense and distance-improving:
$r^{\mathrm{team}}_{k}\;=\;\begin{cases}1,&\text{if at goal at }k,\\[4.0pt] \max\!\bigl(0,\ (d_{k-1}-d_{k})/d_{0}\bigr),&\text{otherwise}.\end{cases}$
We set the team–local mixing coefficient to a fixed number
$\lambda_{\mathrm{plan}}=0.50.$
For each agent $i\in\{\mathrm{Planner},\mathrm{Tool}\}$ with mask $m^{i}_{k}\in\{0,1\}$,
$r^{i}_{k}\;=\;\lambda_{\mathrm{plan}}\,r^{\mathrm{team}}_{k}\;+\;(1-\lambda_{\mathrm{plan}})\,m^{i}_{k}\,r^{i,\mathrm{loc}}_{k}.$
Local rewards are convex combinations $r^{i,\mathrm{loc}}_{k}=\sum_{\ell}c^{i}_{\ell}\,s^{i}_{\ell,k}$ with fixed $\sum_{\ell}c^{i}_{\ell}=1$.
@@ A.1.4 Plan-Path reward design
$c^{\mathrm{Planner}}_{\mathrm{fmt}}=0.20,\quad c^{\mathrm{Planner}}_{\mathrm{leg}}=0.40,\quad c^{\mathrm{Planner}}_{\mathrm{sp}}=0.40.$
Component scores at turn $k$ (action $a_{k}\in\{\mathrm{U,D,L,R}\}$; $\mathcal{N}$ denotes passable neighbors; SPNext is 1 if $a_{k}$ lies on at least one shortest path from $s_{k-1}$ to goal, else 0):
$s^{\mathrm{Planner}}_{\mathrm{fmt},k}=\mathbf{1}\{a_{k}\in\{\mathrm{U,D,L,R}\}\},$
$s^{\mathrm{Planner}}_{\mathrm{leg},k}=\mathbf{1}\{\text{next cell in-bounds and not a wall}\},$
$s^{\mathrm{Planner}}_{\mathrm{sp},k}=\begin{cases}1,&\text{if }\textsc{SPNext}(a_{k}){=}1,\\ 0,&\text{otherwise}.\end{cases}$
@@
$m^{\mathrm{Planner}}_{k}=\mathbf{1}\{\text{map known and shortest-path oracle available at }k\}.$
@@
$c^{\mathrm{Tool}}_{\mathrm{fmt}}=0.10,\quad c^{\mathrm{Tool}}_{\mathrm{exec}}=0.40,\quad c^{\mathrm{Tool}}_{\mathrm{shape}}=0.50.$
Component scores (let $\phi_{k}=-d_{k}$ be the potential used in shaping):
$s^{\mathrm{Tool}}_{\mathrm{fmt},k}=\mathbf{1}\{\text{action list parsable as }[\texttt{"U","D","L","R"}]\},$
$s^{\mathrm{Tool}}_{\mathrm{exec},k}=\mathbf{1}\{\text{no invalid move applied; simulation advances}\},$
$s^{\mathrm{Tool}}_{\mathrm{shape},k}=\mathbf{1}\{\phi_{k}\geq\phi_{k-1}\},$
i.e., the potential does not decrease. Mask:
$m^{\mathrm{Tool}}_{k}=\mathbf{1}\{\text{execution logs and potentials }(\phi_{k-1},\phi_{k})\text{ available}\}.$
A.1.5 Sokoban reward design
We consider Sokoban with horizon $K$ on a fixed grid. Let $B$ be the number of boxes and $b_{k}$ the number of boxes on goal at turn $k$. Team reward is dense in box-on-goal ratio with terminal success at completion:
$r^{\mathrm{team}}_{k}\;=\;\begin{cases}1,&\text{if all boxes on goals at }k,\\[4.0pt] b_{k}/B,&\text{otherwise}.\end{cases}$
We set the team–local mixing coefficient to a fixed number
$\lambda_{\mathrm{sok}}=0.40.$
For each agent $i\in\{\mathrm{Planner},\mathrm{Tool}\}$ with mask $m^{i}_{k}\in\{0,1\}$,
$r^{i}_{k}\;=\;\lambda_{\mathrm{sok}}\,r^{\mathrm{team}}_{k}\;+\;(1-\lambda_{\mathrm{sok}})\,m^{i}_{k}\,r^{i,\mathrm{loc}}_{k}.$
Local rewards are convex combinations $r^{i,\mathrm{loc}}_{k}=\sum_{\ell}c^{i}_{\ell}\,s^{i}_{\ell,k}$ with fixed $\sum_{\ell}c^{i}_{\ell}=1$.
Planner local design.
Coefficients (fixed):
$c^{\mathrm{Planner}}_{\mathrm{fmt}}=0.10,\quad c^{\mathrm{Planner}}_{\mathrm{leg}}=0.45,\quad c^{\mathrm{Planner}}_{\mathrm{dlk}}=0.45.$
Component scores at turn $k$ (action $a_{k}\in\{\mathrm{U,D,L,R}\}$; $\textsc{PushOK}=1$ if a planned push does not collide and stays in-bounds; $\textsc{DeadlockFree}=1$ if the move avoids standard static corner deadlocks for boxes not on goals):
$s^{\mathrm{Planner}}_{\mathrm{fmt},k}=\mathbf{1}\{a_{k}\in\{\mathrm{U,D,L,R}\}\},$
[…]
$s^{\mathrm{Planner}}_{\mathrm{dlk},k}=\begin{cases}1,&\text{if }\textsc{DeadlockFree}=1,\\ 0,&\text{otherwise}.\end{cases}$
@@ A.1.5 Sokoban reward design
$m^{\mathrm{Planner}}_{k}=\mathbf{1}\{\text{grid known and deadlock heuristics evaluable at }k\}.$
Tool (executor/simulator) local design.
Coefficients (fixed):
$c^{\mathrm{Tool}}_{\mathrm{fmt}}=0.10,\quad c^{\mathrm{Tool}}_{\mathrm{exec}}=0.30,\quad c^{\mathrm{Tool}}_{\mathrm{pot}}=0.60.$
Let $\psi_{k}=-\sum_{x\in\text{boxes}}\min_{g\in\text{goals}}\bigl(|x_{r}-g_{r}|+|x_{c}-g_{c}|\bigr)$ be the box-to-goal potential (larger is better). Component scores:
$s^{\mathrm{Tool}}_{\mathrm{fmt},k}=\mathbf{1}\{\text{action list parsable; symbols match }\{\mathrm{U,D,L,R}\}\},$
$s^{\mathrm{Tool}}_{\mathrm{exec},k}=\mathbf{1}\{\text{no illegal push; no wall/box collision}\},$
$s^{\mathrm{Tool}}_{\mathrm{pot},k}=\mathbf{1}\{\psi_{k}\geq\psi_{k-1}\}.$
Mask:
$m^{\mathrm{Tool}}_{k}=\mathbf{1}\{\text{execution logs and potentials }(\psi_{k-1},\psi_{k})\text{ available}\}.$
A.2 Experiment Details
A.2.1 Training Details

### v4

@@ B.2 Code reward design
$c^{\mathrm{Tester}}_{\mathrm{valid}}=0.20,\quad c^{\mathrm{Tester}}_{\mathrm{nr}}=0.80.$
$c^{\mathrm{Tester}}_{\mathrm{cov}}=0.80.$
@@
$\cancel{s^{\mathrm{Tester}}_{\mathrm{cov}}=\begin{cases}\min\!\left(1,\ \dfrac{\mathrm{MutScore}}{\tau_{\mathrm{mut}}}\right),&\text{mutation analysis available},\\[6.0pt] 0,&\text{otherwise},\end{cases}}$
$s^{\mathrm{Tester}}_{\mathrm{nr},t}=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\mathbf{1}\{\textsc{Run}(u,\text{code}^{\star})=\textsf{pass}>\tau_{mut}\},$
where $\tau_{\mathrm{mut}}=0.60$.
Mask:
$m^{\mathrm{Tester}}_{t}=\mathbf{1}\{\text{test runner and mutation/coverage reports available at }t\}.$
The Tester local reward $r^{\mathrm{loc}}_{t,\mathrm{Tester}}$ is obtained by combining $c^{\mathrm{Tester}}_{\cdot}$, $s^{\mathrm{Tester}}_{\cdot,t}$ and $m^{\mathrm{Tester}}_{t}$ via the generic local-reward formula.
B.3 Sudoku reward design
We consider $N{\times}N$ Sudoku. Let $h_{t}$ be the answer action at turn $t$ and $\textsc{Solved}(\cdot)$ check row/column/subgrid validity. Team reward is a sparse success signal at termination, broadcast across turns:
$r^{\mathrm{team}}_{t}\;=\;\mathbf{1}\{\ \textsc{Solved}(h){=}\textsf{true}\}\in\{0,1\},\qquad\forall t.$
Local rewards.
Each agent $i$ at turn $t$ uses a masked convex combination of verifiable sub-scores $s^{i}_{\ell,t}\in[0,1]$:
$r^{\mathrm{loc}}_{t,i}\;=\;m_{t,i}\sum_{\ell}c^{i}_{\ell}\,s^{i}_{\ell,t},\qquad\sum_{\ell}c^{i}_{\ell}=1.$
Reasoner local design.
Coefficients :
$c^{\mathrm{Reasoner}}_{\mathrm{fmt}}=0.1,\quad c^{\mathrm{Reasoner}}_{\mathrm{legal}}=0.1,\quad c^{\mathrm{Reasoner}}_{\mathrm{prog}}=0.80.$
Component scores (let $G_{t}$ be the current grid, $G_{t-1}$ the previous grid; 0 denotes empty):
@@ B.3 Sudoku reward design
$s^{\mathrm{Reasoner}}_{\mathrm{legal},t}=\mathbf{1}\{\text{no row/column/subgrid duplicates in }G_{t}\},$
$s^{\mathrm{Reasoner}}_{\mathrm{prog},t}=\frac{1}{N^{2}}\sum_{r,c}\mathbf{1}\{G_{t-1}[r,c]{=}0,\ G_{t}[r,c]{\neq}0\}.$
@@
$c^{\mathrm{Tool}}_{\mathrm{fmt}}=0.10,\quad c^{\mathrm{Tool}}_{\mathrm{exec}}=0.10,\quad c^{\mathrm{Tool}}_{\mathrm{san}}=0.80.$
@@
$s^{\mathrm{Tool}}_{\mathrm{fmt},t}=\mathbf{1}\{\text{API/schema valid; values in }[1,N];\ \text{indices in bounds}\},$
$s^{\mathrm{Tool}}_{\mathrm{exec},t}=\mathbf{1}\{\text{no runtime error/timeout when applying edits}\},$
$s^{\mathrm{Tool}}_{\mathrm{san},t}=\begin{cases}1,&\text{if all applied edits satisfy local Sudoku constraints},\\ 0,&\text{otherwise}.\end{cases}$
@@
B.4 Plan-Path reward design
We consider 2D grid path planning on a $H{\times}W$ map with horizon $T$ and four-neighborhood moves. Let $d_{t}$ be the Manhattan distance from the current position to the goal at turn $t$ and $d_{0}=\max(1,\text{initial distance})$ for normalization. Team reward is dense and distance-improving:
$r^{\mathrm{team}}_{t}\;=\;\begin{cases}1,&\text{if at goal at }t,\\[4.0pt] \max\!\bigl(0,\ (d_{t-1}-d_{t})/d_{0}\bigr),&\text{otherwise}.\end{cases}$
Local rewards are masked convex combinations
$r^{\mathrm{loc}}_{t,i}\;=\;m_{t,i}\sum_{\ell}c^{i}_{\ell}\,s^{i}_{\ell,t},\qquad\sum_{\ell}c^{i}_{\ell}=1.$
@@ B.4 Plan-Path reward design
$c^{\mathrm{Planner}}_{\mathrm{fmt}}=0.10,\quad c^{\mathrm{Planner}}_{\mathrm{leg}}=0.10,\quad c^{\mathrm{Planner}}_{\mathrm{sp}}=0.80.$
Component scores at turn $t$ (action $a_{t}\in\{\mathrm{U,D,L,R}\}$; $\mathcal{N}$ denotes passable neighbors; SPNext is 1 if $a_{t}$ lies on at least one shortest path from $s_{t-1}$ to goal, else 0):
$s^{\mathrm{Planner}}_{\mathrm{fmt},t}=\mathbf{1}\{a_{t}\in\{\mathrm{U,D,L,R}\}\},$
$s^{\mathrm{Planner}}_{\mathrm{leg},t}=\mathbf{1}\{\text{next cell in-bounds and not a wall}\},$
$s^{\mathrm{Planner}}_{\mathrm{sp},t}=\begin{cases}1,&\text{if }\textsc{SPNext}(a_{t}){=}1,\\ 0,&\text{otherwise}.\end{cases}$
@@
$m^{\mathrm{Planner}}_{t}=\mathbf{1}\{\text{map known and shortest-path oracle available at }t\}.$
Tool (executor/simulator) local design.
Coefficients:
$c^{\mathrm{Tool}}_{\mathrm{fmt}}=0.10,\quad c^{\mathrm{Tool}}_{\mathrm{exec}}=0.10,\quad c^{\mathrm{Tool}}_{\mathrm{shape}}=0.80.$
Component scores (let $\phi_{t}=-d_{t}$ be the potential used in shaping):
$s^{\mathrm{Tool}}_{\mathrm{fmt},t}=\mathbf{1}\{\text{action list parsable as }[\texttt{``U",``D",``L",``R"}]\},$
$s^{\mathrm{Tool}}_{\mathrm{exec},t}=\mathbf{1}\{\text{no invalid move applied; simulation advances}\},$
$s^{\mathrm{Tool}}_{\mathrm{shape},t}=\mathbf{1}\{\phi_{t}\geq\phi_{t-1}\},$
i.e., the potential does not decrease. Mask:
$m^{\mathrm{Tool}}_{t}=\mathbf{1}\{\text{execution logs and potentials }(\phi_{t-1},\phi_{t})\text{ available}\}.$
B.5 Sokoban reward design
We consider Sokoban with horizon $T$ on a fixed grid. Let $B$ be the number of boxes and $b_{t}$ the number of boxes on goal at turn $t$. Team reward is dense in box-on-goal ratio with terminal success at completion:
$r^{\mathrm{team}}_{t}\;=\;\begin{cases}1,&\text{if all boxes on goals at }t,\\[4.0pt] b_{t}/B,&\text{otherwise}.\end{cases}$
Local rewards are masked convex combinations
$r^{\mathrm{loc}}_{t,i}\;=\;m_{t,i}\sum_{\ell}c^{i}_{\ell}\,s^{i}_{\ell,t},\qquad\sum_{\ell}c^{i}_{\ell}=1.$
Planner local design.
Coefficients (fixed):
$c^{\mathrm{Planner}}_{\mathrm{fmt}}=0.10,\quad c^{\mathrm{Planner}}_{\mathrm{leg}}=0.10,\quad c^{\mathrm{Planner}}_{\mathrm{dlk}}=0.80.$
Component scores at turn $t$ (action $a_{t}\in\{\mathrm{U,D,L,R}\}$; $\textsc{PushOK}=1$ if a planned push does not collide and stays in-bounds; $\textsc{DeadlockFree}=1$ if the move avoids standard static corner deadlocks for boxes not on goals):
$s^{\mathrm{Planner}}_{\mathrm{fmt},t}=\mathbf{1}\{a_{t}\in\{\mathrm{U,D,L,R}\}\},$
[…]
$s^{\mathrm{Planner}}_{\mathrm{dlk},t}=\begin{cases}1,&\text{if }\textsc{DeadlockFree}=1,\\ 0,&\text{otherwise}.\end{cases}$
Mask:
$m^{\mathrm{Planner}}_{t}=\mathbf{1}\{\text{grid known and deadlock heuristics evaluable at }t\}.$
@@ B.5 Sokoban reward design
$c^{\mathrm{Tool}}_{\mathrm{fmt}}=0.10,\quad c^{\mathrm{Tool}}_{\mathrm{exec}}=0.10,\quad c^{\mathrm{Tool}}_{\mathrm{pot}}=0.80.$
Let $\psi_{t}=-\sum_{x\in\text{boxes}}\min_{g\in\text{goals}}\bigl(|x_{r}-g_{r}|+|x_{c}-g_{c}|\bigr)$ be the box-to-goal potential (larger is better). Component scores:
$s^{\mathrm{Tool}}_{\mathrm{fmt},t}=\mathbf{1}\{\text{action list parsable; symbols match }\{\mathrm{U,D,L,R}\}\},$
$s^{\mathrm{Tool}}_{\mathrm{exec},t}=\mathbf{1}\{\text{no illegal push; no wall/box collision}\},$
$s^{\mathrm{Tool}}_{\mathrm{pot},t}=\mathbf{1}\{\psi_{t}\geq\psi_{t-1}\}.$
@@
$m^{\mathrm{Tool}}_{t}=\mathbf{1}\{\text{execution logs and potentials }(\psi_{t-1},\psi_{t})\text{ available}\}.$
B.6 Outcome-only reward design
The shaped rewards in Sections B 1–5 provide rich, task-specific feedback (e.g., shortest-path signals in Plan-Path and deadlock heuristics in Sokoban). To isolate the contribution of such complex shaping from that of the AT-GRPO algorithm itself, we additionally consider a simplified outcome-only reward design used in our ablation studies.
The team reward is strictly binary and episodic. Let $\mathbb{I}(\text{Success})$ denote the environment success indicator. The team reward is defined and broadcast over turns as
$r^{\mathrm{team}}_{t}\;=\;\mathbb{I}(\text{Success}),\qquad\forall t.$
The per-agent local reward in this setting is an auxiliary binary signal that only checks whether agent $i$ produced a validly formatted action (e.g., correct API call or JSON structure). Let $\mathbb{I}(\text{FmtValid}^{i}_{t})$ denote the indicator that the output of agent $i$ at turn $t$ satisfies all formatting constraints. We define
$r^{\mathrm{loc}}_{t,i}\;=\;r^{i,\mathrm{out}}_{t}\;=\;\mathbb{I}\big(\text{FmtValid}^{i}_{t}\big).$
The final per-agent reward $r_{t,i}$ is then combined according to Eq. 3,
$r_{t,i}\;=\;\alpha\,r^{\mathrm{team}}_{t}\;+\;r^{\mathrm{loc}}_{t,i},$
where we use a fixed, task-independent $\alpha$ shared with the shaped-reward configurations. By utilizing sparse episodic rewards and simple formatting checks rather than dense shaping signals, this outcome-only configuration is significantly more general and provides a baseline for our algorithm ablation.
B.7 Theoretical Justification for Greedy Turn-Level Transitions
In this section, we formally justify the optimality of greedy selection based on environment-verified rewards. Consider the underlying MDP with optimal action-value function $Q^{*}(s,a)$. The Bellman optimality principle implies that any policy $\pi^{*}$ satisfying $\pi^{*}(s)\in\arg\max_{a}Q^{*}(s,a)$ is optimal [sutton2018rl].
We operate in a setting where the environment returns an outcome-based verifiable reward, $r_{\mathrm{ver}}(s,a)$, for each action. We posit that this reward acts as a monotonic proxy for the true value function $Q^{*}(s,a)$: a higher verification score directly corresponds to a higher probability of final success. Consequently, maximizing the immediate verifiable reward is structurally equivalent to maximizing the long-term optimal value. We formalize this alignment as follows:
Assumption 1 (Monotonicity of Verification Feedback).
For any state $s$ and actions $a_{1},a_{2}$, the verifiable reward preserves the ordering of the optimal action-value function:
$r_{\mathrm{ver}}(s,a_{1})>r_{\mathrm{ver}}(s,a_{2})\implies Q^{*}(s,a_{1})\geq Q^{*}(s,a_{2}).$ (4)
This implies that $r_{\mathrm{ver}}(s,\cdot)$ and $Q^{*}(s,\cdot)$ induce consistent rankings over the action space at any state $s$.
Lemma 1 (Equivalence of Maximizers).
Under Assumption 1, the set of actions maximizing the verifiable reward is a subset of the actions maximizing the optimal $Q$-function:
$\arg\max_{a}r_{\mathrm{ver}}(s,a)\subseteq\arg\max_{a}Q^{*}(s,a).$ (5)
Proof.
Let $a^{*}_{\mathrm{ver}}\in\arg\max_{a}r_{\mathrm{ver}}(s,a)$. Suppose, for the sake of contradiction, that there exists an action $a^{\prime}$ such that $Q^{*}(s,a^{\prime})>Q^{*}(s,a^{*}_{\mathrm{ver}})$. By the contrapositive of Assumption 1, strict inequality in $Q^{*}$ implies strict inequality in $r_{\mathrm{ver}}$ (given consistent rankings), which would imply $r_{\mathrm{ver}}(s,a^{\prime})\geq r_{\mathrm{ver}}(s,a^{*}_{\mathrm{ver}})$. Since $a^{*}_{\mathrm{ver}}$ is a maximizer, strict inequality is impossible. If equality holds, $a^{\prime}$ is also a maximizer of $r_{\mathrm{ver}}$, and by the consistency assumption, it must share the same optimal $Q$-value. Thus, any action maximizing $r_{\mathrm{ver}}$ necessarily maximizes $Q^{*}$. ∎
Proposition 1 (Optimality of Verifier-Greedy Policy).
Let $\pi_{\mathrm{ver}}$ be a deterministic policy such that $\pi_{\mathrm{ver}}(s)\in\arg\max_{a}r_{\mathrm{ver}}(s,a)$ for all states $s$. Under Assumption 1, $\pi_{\mathrm{ver}}$ is an optimal policy.
Proof.
By Lemma 1, selecting an action that maximizes the immediate verification score ensures that $\pi_{\mathrm{ver}}(s)\in\arg\max_{a}Q^{*}(s,a)$. Consequently, $\pi_{\mathrm{ver}}$ satisfies the Bellman optimality equation at every state. ∎
In our implementation, we approximate this policy by sampling candidate actions and greedily selecting the one with the highest $r_{\mathrm{ver}}$. Proposition 1 guarantees that this strategy effectively performs a greedy search over the support of sampled actions with respect to the true optimal value function $Q^{*}$, avoiding the myopic bias typically associated with greedy transitions.
Appendix C Experiment Details
C.1 Training Details